write in the form of x=g(x)
e^x+arcsin(x)=0
1
Expert's answer
2020-11-09T19:44:32-0500

"e^x + arcsin(x) =0\\
e^x =-arcsin(x)\\
-e^x=arcsin(x)\\
sin(-e^x)=x\\
x=sin(-e^x)\\
x=-sin(e^x)\\
\textsf{This is in the form x=g(x)}\\
\textsf{Where } g(x)=-sin(e^x)"